Two identical closely spaced circular disks form a parallel-plate capacitor. Transferring 1.9×10^9 electrons from one disk to the other causes the electric field strength between them to be 1.1×10^5 N/C . What are the diameters of the disks?
Answer #1

Solution)

Given, Electric field E = q / ε A = 1.1*10^5 N/C

Charge transferred

q = n * e

Here, e = 1.6 * 10-19 C

where q = charge on plates = 1.9 * 10^ 9 * 1.6 * 10 -19C

We know, ε = permitivity=8.85 * 10 -12 C2 / N m 2

  

A= area of cross section = πd 2 /4

So, E = 4q / [ ε π d^ 2]

Substituting values, we get

4 * 1.9 * 10^ 9 * 1.6 * 10 -19 / (8.85 * 10^-12 * 3.14 * d^2 ) = 1.1* 10^5   

4 * 1.9 * 10^9 * 1.6 * 10 ^-19 = 1.1* 10^5 * 8.85 * 10^-12 * 3.14 * d^2

d= 0.0199 m (Ans)

======

Good luck!:)